A company estimates that it will need $71,000 in 14 years to replace a computer. If it establishes a sinking fund by making fixed monthly payments into an account paying 4.2% compounded monthly, how much should each payment be?

The amount of each payment should be$

(Round to the nearest cent.)
Answer #1

Monthly installment Formula:

where

E is EMI

P is Principal Loan Amount =71000

r is the rate of interest calculated on a monthly basis = 4.2%

n is loan term/tenure/duration in number of months = 14*12 = 168

So,

E = 71,000*0.042*(1+0.042)168/((1+0.042)168 - 1)

= 2982*1004.1/(1004.1-1)

= $2984.97

Please upvote if you have liked my answer, would be of great help. Thank you.